# GestureGPT: Toward Zero-shot Free-form Hand Gesture Understanding with Large Language Model Agents

### v4

@@ 3.1. Gesture Description Agent
The Gesture Description Agent is crucial for translating video-captured gestures into natural language descriptions that is understandable by LLMs. We choose LLM since it has richer common knowledge and stronger inference than existing LMM that we can access, which are vital to deal with context-ware free-form gesture understanding tasks.
3.1.1   Rule-Based Gesture Description Generation
We design a set of 6 rules to encode the hand gesture in terms of finger flexion, proximity, contact, direction, as well as palm orientation and hand position (Table 1). Each state is calculated based on the coordinates of hand landmarks generated by MediaPipe (Zhang et al., 2020) with videos captured by a RGB camera. For each rule, parameters (such as the distance between fingers to determine their proximity) are trained on third-view public gesture dataset, and tested on first-view and third-view datasets. Test results show an error rate of 2.3% for third-view and 6.3% for first-view samples, indicating the rules’ efficacy in retaining accurate information from gestures. The detailed process of rule definition and training, as well as the evaluation metrics for these rules, are thoroughly described in Appendix A.
@@ 3.1.1   Rule-Based Gesture Description Generation
Rule Name Applicable to Value
Finger Flexion Thumb, Index, Middle, Ring, Pinky 1: Straight;
@@
Finger Proximity Index-Middle, Middle-Ring, Ring-Pinky 1: Pressed Together;
@@
Thumb Fingertip Contact Thumb-Index, Thumb-Middle, Thumb-Ring, Thumb-Pinky 1: Contact;
@@
Pointing Direction Thumb 1: Upward;
@@
Palm Orientation Palm One-hot encoding: [Left, Right, Down, Up, Inward, Outward];
@@
Hand Position Hand Float Coordinates
@@
The gesture period of interest starts when users raises their hands to or above their chest level. Frames within a gesture period are processed at 0.2-second intervals, with each sampled frame undergoing rule-based calculations. However, the natural language descriptions concatenated from each rule are too long to fit into a prompt of an LLM. So we introduce a gesture state matrix that contains vectors of each frame, which capitalizes on LLMs’ proficiency with code-formatted content (Gao et al., 2023b).
3.1.2   Rules Calculation Method
Flexion of a finger is calculated as the total bending angle of each joint. For thumb it is the bending angle of the ip joint, and for other fingers, it is the bending angle of the pip and dip joint. Then, two parameters $straight\_threshold$ and $bent\_threshold$ are set to determine if the finger is straight, bent, or ‘unsure’ if the result falls between them. Since thumb has a different joint structure compared with other fingers, a new pair of thresholds are specially set for thumb.
Proximity of two fingers A and B is calculated as the average minimal distance from each finger’s joint to the other finger. Two thresholds i.e. $together\_threshold$ and $separated\_threshold$ are set to determine if the two fingers are pressed together, separated, or ‘unsure’ if the result falls between them.
Contact of thumb and another finger is computed as the distance between their fingertips. Then, two thresholds, i.e. $contact\_threshold$ and $not\_contact\_threshold$ are set to determine if the two fingers’ fingertips are in contact or not, or ‘unsure’ if the result falls between them.
Point direction of thumb is computed as the direction from thumb’s mcp joint to tip joint. Then, it is compared with two reference vectors representing upward and downward. If a reference vector has the minimal angle with the palm orientation vector and the angle is below $angle\_threshold$, the reference vector would be the thumb’s pointing direction. Note that it is only applicable when thumb is straight. If thumb is bent, the result is set to ‘unsure’. This is especially useful when discriminating between gestures like ‘thumb up’ and ‘thumb down’.
Palm orientation is computed as the direction to which the palm is facing. It is computed by the cross product of two vectors within the plane of the palm. (Fig 4). Then, the direction is compared with six reference vectors representing upward, downward, left, right, inward and outward. If a reference vector has the minimal angle with the palm orientation vector and the angle is below $angle\_threshold$, the reference vector would be the palm orientation. Otherwise the orientation of palm is set as ‘unsure’.
The pseudocode for the rules are given in Appendix A.2., The rule-based module evaluation and the matrix can be found in Appendix A.3.
3.1.3   Gesture Summary Description Generation
@@
The system needs to extract the rich information embedded in the matrix and generate a concise summary of gesture descriptions to pass to the Gesture Inference Agent. A significant challenge encountered by the agent is the novelty of the state matrix data type; the pre-trained data of the LLM likely does not contain exactly similar data. We employ a chain-of-thought (Wei et al., 2023) process to guide the LLM step-by-step, supplemented by high-quality expert examples that demonstrate how the analysis should be conducted. We choose to use two prompts to generate the hand pose (finger flexion, finger proximity, fingertip contact, thumb direction and palm orientation) and hand movement description separately. As is shown in Figure 3, the Gesture Description Agent first converts pose-related rows of the matrix to descriptions, along with the time span during which this gesture happened. The time span is then used to find the corresponding part of movement-related rows, and then another prompt is used to convert it to movement description. As long as the agent correctly finds the pose movement, irrelevant movements are naturally filtered out.
@@
For pose description, we developed a prompt structured into three parts shown in Figure 5: introduction, procedure, examples. The introduction outlines the task, describes the input matrix, and explains interactive gestures to distinguish between types of gestures and understand their life cycle. The procedure section teaches the agent to decompose the matrix, encouraging a human-like process of understanding and summarizing data, while suggesting transcription to counter LLM’s forgetfulness and improve accuracy. The last part provides examples of static and dynamic gestures and a structured prompt to enhance LLM performance by leveraging its knowledge base for complex task execution. The well-structured prompt ensures the LLM model correctly processes the matrix data type. The hand movement description prompt shares a similar structure.
Figure 5. Formatted Prompt For Gesture Description Agent.
@@ 3.2. Gesture Inference Agent
Upon receiving gesture descriptions from Gesture Description Agent, Gesture Inference Agent engages in dialogue with Context Management Agent. This conversational exchange is pivotal for the Gesture Inference Agent to discern the user’s actual intent in the context, i.e., associating the gesture with a target system function from a list of potential functions supplied by Context Management Agent. This setup epitomizes a collaborative conversation aimed at overcoming the challenges of gesture ambiguities under different interactive scenarios and context.
@@
Figure 6. Formatted Prompt For Gesture Inference Agent.
The introduction part includes a Task Introduction, a Gesture Description Introduction, and a Context Management Agent Introduction (its conversational counterpart). A critical aspect to convey in this part is that the initial gesture description may contain errors. Therefore, the Gesture Inference Agent must utilize context information to help with the inference, correct potential mistakes, and make decisions.
Behavioral guidelines for the agent consist of seven directives and five prohibitions, acquired by iteratively updating the prompts during implementation. When new error tendencies are observed, corrections are introduced into the prompt, akin to a teacher’s guidance. Then, we define the output format. We opt for JSON due to its ease of parsing, allowing for efficient extraction of useful information for subsequent conversational rounds. Importantly, we require the agent to first articulate its thoughts before posing questions or drawing conclusions. This approach has been proven to make the agent’s behavior more reasonable. Once the agent deems that it is confident to decide or that no further context can be gleaned, it proceeds to make a final decision, listing the top-5 possible functions for the current interface, ranked from most to least likely. The Gesture Inference Agent emphasizes the importance of precise communication, contextual reasoning, and analytical capabilities in interpreting and responding to gesture-based interactions.
@@ 3.3. Context Management Agent
Context management is fundamental for several reasons. First, it facilitates fast interpretation of and dynamic adaptation to diverse types of context information, empowering the system to make informed decisions beyond the limitations of predefined rules. This adaptability significantly enhances system responsiveness and allows for a more precise tailoring of responses to user gestures. Additionally, it streamlines the handling of complex contextual interactions, making the system more accessible to both users, developers and agents.
The overall prompt is structured into two parts shown in Figure 7.
Figure 7. Formatted Prompt For Context Management Agent.
The introduction part includes a Task Introduction and a Context Library Introduction. The Context Library Introduction introduces the currently accessible contexts and is designed to be adjustable. Similar to the gesture inference agent, iterative implementation guides the Context Management Agent’s requirements and prohibitions. There are 4 requirements and 2 prohibitions. The output is structured in JSON format for ease of parsing, allowing for efficient extraction of useful information and facilitation of subsequent conversation rounds.
3.3.1   Context Library Operations
@@ 4. Evaluation
We selected four contexts in our experiments to evaluate GestureGPT performance in different context setting:
@@
  • Interaction History: Insights into the user’s recent interactions.
@@
We informed participants that both video and eye movement data would be collected during the study, and we assured them of the confidentiality and safety of their data. The study is IRB approved by our local institution. All participants provided informed consent. In both experiments, we asked participants to perform the gesture as they see fit to finish the task. Both static hand poses and dynamic gestures involving hand movements are permitted.
We employed the OpenAI GPT-4 model as the underlying architecture for our triple-agent system. Specifically, we utilized the gpt-4-1106-preview version for our evaluations. We configured the request temperature to 0 so that the model’s output is as consistent as possible. Apart from this, we adhered to OpenAI’s default settings for other parameters. To mitigate the effects of randomness and enhance the reliability of our findings, we run the experiment repeatedly for three times. The aggregated results from these iterations were used to substantiate our conclusions.
@@ 4.1. Experiment 1: Augmented Reality-Based Smart Home IoT Control
4.1.1   Experiment Setting and Procedure
  • Experimental Platform and Data Collection - The Microsoft HoloLens 2 served as the primary experimental platform, offering APIs to capture user gaze and hand gesture data accurately. The experimental environment was developed using Unity (version 2020.3.24f1), the Mixed Reality Toolkit (MRTK 2.8.0), and the OpenXR Plugin (1.7.0). The devices were represented with 3D models anchored in the space: a light, a smart cabinet, a smart screen, an oven, and an air cleaner.
@@ 4.1.1   Experiment Setting and Procedure
    • Gaze Data: User gaze data was captured using the HoloLens 2 Gaze API and saved as 3D spatial coordinates.
@@
    • External Context: There might be context information that is external to our system, which can significantly impact the grounding reasoning. We defined several external contexts corresponding to different tasks to understand if our system can correctly leverage those.
  • Task Descriptions - Specific eight tasks assigned to participants related to smart device control.
  • Participants - We recruited 16 participants from three local schools, compensating them at a rate of \$12 per hour. Their ages ranged from $15$ to $35$ years ($\text{MEAN}=26.625,\text{SD}=5.325$), comprising $13$ males and $3$ females.
@@
Detailed information about the experiment simulation interface, the list of devices and their functions, and the task list is provided in Appendix B.1.
@@ 4.1.2   Results Analysis
We run the collected data through our system with four different context settings to evaluate GestureGPT’s performance, respectively: 1) the baseline without any context other than the function list(Baseline), 2) with gaze information (Only Gaze), 3) with interaction history and external information (Only History and External), and 4) all contexts are available (All). We also provide the results of Random Guess for comparison, which randomly selects from all candidate functions. We run the results three times under each setting to provide robust conclusions. The main results are listed in the left part of Table 2. We also provide a corresponding illustration in Figure 8.
@@
When comparing the specific benefits between gaze and history and external information, we empirically find that gaze provides much more significant benefits than history and external information. As can be seen from Table 2, the Only Gaze setting achieves 35.16%/83.59% at Top-1/Top-5 accuracy, outperforming the 23.18%/49.48% achieved by the Only History and External setting by a considerable margin. We attribute this to the semantic similarity of candidate functions shared by different home devices as well as their spatial distribution across the entire room space. These two factors significantly increase the reasoning difficulty under the Smart Home scenario. Facing such challenging situations, GestureGPT can fully exploit the gaze information as well as its own commonsense knowledge to finally locate and zoom in on possible candidates through multi-step, collaborative spatial reasoning. As explained in Section 3.3.1, when the Gesture Inference Agent requests gaze information, the Context Management Agent not only outputs the gaze coordinates but also identifies the relevant device within the gaze path using external tool-augmented Python scripts.
@@
Figure 8. Illustration of results from Experiment 1: Home Scenario, encompassing all 8 tasks (across the x-axis) and 4 settings (indicated by hue). For each task, the 4 settings are arranged from left to right as follows: Baseline, Only Gaze, Only History and External, and All. Each specific bar is further divided into Top 1 / Top 3 / Top 5 accuracy (Top 1, Top 3, Top 5 in gradient brown) and Negative outcomes (in cyan), totaling 48 data points.
@@
Moreover, while integrating All context setting achieves the highest Top-1 performance, it unexpectedly underperforms in the Top-5 metric compared to the Gaze Only setting (79.69% vs 83.59%). Through further case examination, we hypothesize that an excess of context can sometimes disrupt the agent’s analysis, leading to the following behaviors: 1) Irrelevant context information can at times mislead the agent, for example, the mention of ”fingerprint unlocking”(in task 1) within external information might lead the agent to infer fingerprint recognition as the unlocking method, rather than a gesture. 2) An abundance of context may cause the agent to overlook certain information, for example, in the ’Open the air purifier’ task (in task 5) where the agent incorrectly assumes the device is on based on the external information ’The air purifier’s sensor detected that the current environment has heavy cooking fumes’, disregarding its actual OFF status. These insights will be further discussed in Section 5.3 regarding optimal context selection.
@@
Finally, we observe that Tasks 2, 3, and 7 exhibit relatively high baseline performance, all of which involve operations related to ”switching”. For example, in Task 2 ’Increasing Light Brightness’, GestureGPT first recognizes that the gesture involves significant ’movement’, leading the agent to consider that the gesture may be related to some type of value control function. Then, by considering the current status of devices—specifically that the lights were on while other devices with similar switch and control functionalities were mostly off—the agent was able to infer that the gesture is related to Light device control. Similarly, for Tasks 3 and 7, GestureGPT employed a comparable analytical approach to discern the correct answer. This illustrates GestureGPT’s strength in analyzing gestures in conjunction with the current states of devices, thereby accurately interpreting user intentions even in the baseline context.

### v5

@@ 3.1. Gesture Description Agent
The Gesture Description Agent is crucial for translating video-captured gestures into natural language descriptions that is understandable by LLMs. We choose LLM since it has richer common knowledge and stronger inference than existing Large Multimodal Models (LMMs) that we can access, which is vital to deal with context-aware free-form gesture understanding tasks.
3.1.1. Rule-Based Gesture Description Generation
We design a set of 6 rules to encode the hand gesture in terms of finger flexion, proximity, contact, direction, as well as palm orientation and hand position (Table 1). The detailed rule definition can be found in Appendix A.1.
Each state is calculated using the coordinates of hand landmarks, as shown in Fig 4, which are generated by MediaPipe (Zhang et al., 2020) from videos captured with an RGB camera. For each rule, parameters (such as the distance between fingers to determine their proximity) are trained on a third-person view (i.e., viewed from an external perspective, typically through a camera not aligned with the subject’s viewpoint) public gesture dataset, and tested on first-person view (i.e., viewed from the subject’s own perspective, typically through a head-mounted camera) and third-person view datasets. Test results show an error rate of 2.3% for third-person view and 6.3% for first-person view samples, indicating the rules’ efficacy in retaining accurate information from gestures across different viewpoints, validating it a universal solution applicable to various gesture capture devices. The details of parameter training and the evaluation metrics for these rules are thoroughly described in Appendix A. The pseudocode for the rules is provided in Appendix A.3. An evaluation and error analysis of the rules can be found in Appendix A.2.
@@ 3.1.1. Rule-Based Gesture Description Generation
Rule Name Applicable to Calculation Value
Finger Flexion Thumb, Index, Middle, Ring, Pinky measured as the total bending angle of each joint: for the thumb, the IP joint; for other fingers, the PIP and DIP joints. 1: Straight;
@@
Finger Proximity Index-Middle, Middle-Ring, Ring-Pinky calculated as the average minimal distance from each finger’s joint to the other finger. 1: Pressed Together;
@@
Thumb Fingertip Contact Thumb-Index, Thumb-Middle, Thumb-Ring, Thumb-Pinky computed as the distance between their fingertips. 1: Contact;
@@
Pointing Direction Thumb the direction from thumb’s mcp joint to tip joint. 1: Upward;
@@
Palm Orientation Palm computed as the direction to which the palm is facing. One-hot encoding: [Left, Right, Down, Up, Inward, Outward];
@@
Hand Position Hand computed as the geometrical center of a hand by taking average of all 21 landmarks’ coordinates. Float coordinates
@@
The gesture period of interest starts when users raises their hands to or above their chest level. Frames within a gesture period are processed at 0.2-second intervals, with each sampled frame undergoing rule-based calculations. However, the natural language descriptions concatenated from each rule are too long to fit into a prompt of an LLM. So we introduce a gesture state matrix that contains vectors of each frame, which capitalizes on LLMs’ proficiency with code-formatted content (Gao et al., 2023b). The detailed definition of the gesture state matrix is given in Appendix A.4.
3.1.2. Gesture Summary Description Generation
@@
The system needs to extract the rich information embedded in the matrix and generate a concise summary of gesture descriptions for the Gesture Inference Agent. A significant challenge is the gesture state matrix data type; the pre-trained data of the LLM likely does not contain exactly similar data. We employ a chain-of-thought (Wei et al., 2023) process to guide the LLM step-by-step, supplemented by high-quality expert examples that demonstrate how the analysis should be conducted. We choose to use two prompts to generate the hand pose (finger flexion, finger proximity, fingertip contact, thumb direction and palm orientation) and hand movement description separately. As is shown in Figure 3, the Gesture Description Agent first converts pose-related rows of the matrix to descriptions of possible gestures, along with the time span during which this gesture happened. The time span is then used to select the corresponding part of movement-related rows, and another prompt is used to convert it to movement description. As long as the agent correctly finds the valid pose of the gesture, irrelevant movements are naturally filtered out.
@@
For pose description, we developed a prompt structured into three parts shown in Figure 6: Introduction, Procedure and Examples. The introduction outlines the task, describes the the gesture state matrix, and explains the concept of “interactive gestures” versus non-ineractive hand movements, helping the agent to distinguish between different types of gesutres and understand their lifecycle. The procedure section teaches the agent to understand the gesture state matrix and guess possible gestures from the matrix, while addressing some common mistakes to regulate its behaviors. In paticular, it instructs the agent to decompose the matrix, guess possible gestures from each part, and synthesize the conclusions from all parts to get the most possible gestures, encouraging a human-like process of understanding and summarizing data, while using transcription to counter LLM’s forgetfulness and improve accuracy. The last part provides two typical examples of static and dynamic gestures to enhance LLM performance by reflecting back to the guidelines stated before. This well-structured prompt ensures that the LLM model correctly processes the matrix data type. The hand movement description prompt shares a similar structure. The generated description of the gesture will be like:
- Thumb and Index Finger transition from non-specific/bent to straight.
- Middle, Ring, and Pinky Fingers transition from bent to straight.
- Fingers start close together and then spread apart.
- Thumb starts in contact with all fingertips but then moves away.
- Palm orientation starts facing left but becomes non-specific.
- The hand moves left slightly with negligible vertical and depth movements.
Figure 5. Prompt for Gesture Description Agent.
Figure 6. Prompt for Gesture Inference Agent.
@@ 3.2. Gesture Inference Agent
Upon receiving gesture descriptions from Gesture Description Agent, the Gesture Inference Agent engages in a dialogue with Context Management Agent. This conversational exchange is pivotal for the Gesture Inference Agent to discern the user’s actual intent in the context, i.e., associating the gesture with a target system function from a list of potential functions supplied by Context Management Agent. This setup epitomizes a collaborative conversation aimed at overcoming the challenges of gesture ambiguities under different interactive scenarios and context.
@@
The introduction part includes a Task Introduction, a Gesture Description Introduction, and a Context Management Agent Introduction (its conversational counterpart). A critical aspect to convey in this part is that the initial gesture description may contain errors. Therefore, the Gesture Inference Agent must utilize context information to help with the inference, correct potential mistakes, and make decisions. Behavioral guidelines for the agent consist of seven directives and five prohibitions, acquired by iteratively updating the prompts during implementation. When new error tendencies are observed, corrections are introduced into the prompt, akin to a teacher’s guidance. Then, we define the output format. We opt for JSON due to its ease of parsing, allowing for efficient extraction of useful information for subsequent conversational rounds. Importantly, we require the agent to first articulate its thoughts before posing questions or drawing conclusions. This approach has been proven to make the agent’s behavior more reasonable. Once the agent deems that it is confident to decide or that no further context can be gleaned, it proceeds to make a final decision, listing the top-5 possible functions for the current interface, ranked from most to least likely. The Gesture Inference Agent emphasizes the importance of precise communication, contextual reasoning, and analytical capabilities in interpreting and responding to gesture-based interactions.
@@ 3.3. Context Management Agent
A dedicated Context Management Agent represents a significant step forward in the evolution of gesture-based interfaces. By segregating context handling from the inference processes, we streamline system operations and enhance the overall system’s ability to effectively utilize diverse contextual data. This separation allows for more sophisticated and nuanced interaction paradigms, where gestures can be interpreted in varying scenarios with greater precision. Moreover, this architecture facilitates easier updates and scalability, as new contexts or rules can be integrated without disrupting the core inference mechanisms. Consequently, this approach not only addresses previous limitations but also sets a new standard for the development of adaptive, efficient, and user-centric gesture recognition systems. The overall prompt is structured into two parts shown in Figure 7.
Figure 7. Prompt for Context Management Agent.
The introduction includes a Task Introduction and a Context Library Introduction. The latter introduces the currently accessible contexts and is designed to be adjustable. Similar to the Gesture Inference Agent, iterative implementation guides the Context Management Agent’s requirements and prohibitions, with 4 requirements and 2 prohibitions. The output is structured in JSON format for easy parsing, allowing efficient information extraction and facilitating subsequent conversation rounds.
3.3.1. Context Library Operations
@@ 4. Evaluation
We selected four contexts in our experiments to evaluate GestureGPT’s performance in different context settings:
@@
  • Interaction History: A record of the user’s recent interactions.
@@
We informed participants that both video and eye movement data would be collected during the study, and we assured them of the confidentiality and safety of their data. The study is IRB approved by our local institution. All participants provided informed consent. In both experiments, we asked participants to perform the gesture as they see fit to finish the task. Both static hand poses and dynamic gestures involving hand movements were permitted.
We employed the OpenAI GPT-4 model as the underlying architecture for our triple-agent system. Specifically, we utilized the gpt-4-1106-preview version for our evaluations. We configured the request temperature to 0 so that the model’s output is as consistent as possible. Apart from this, we adhered to OpenAI’s default settings for other parameters. To mitigate the effects of randomness and enhance the reliability of our findings, we ran the experiment repeatedly for three times. The aggregated results from these iterations were used to substantiate our conclusions.
@@ 4.1. Experiment 1: Augmented Reality-Based Smart Home IoT Control
4.1.1. Experiment Setting and Procedure
  • Experimental Platform and Data Collection - The Microsoft HoloLens 2 served as the primary experimental platform, offering APIs to capture user gaze and hand gesture data accurately. The experimental environment was developed using Unity (version 2020.3.24f1), the Mixed Reality Toolkit (MRTK 2.8.0), and the OpenXR Plugin (1.7.0). The devices were represented by 3D models anchored in the space: a light, a smart cabinet, a smart screen, an oven, and an air cleaner.
@@ 4.1.1. Experiment Setting and Procedure
    • Gaze Information: User gaze data was captured using the HoloLens 2 Gaze API and saved as 3D spatial coordinates.
@@
    • External Context Information: There might be context information that is external to our system, which can significantly impact the grounding reasoning. We defined several external contexts corresponding to different tasks to understand if our system can correctly leverage those.
  • Task Descriptions - Eight tasks were designed to simulate smart device control (Table 6).
  • Participants - We recruited 16 participants from three local schools, compensating them at a rate of \$12 per hour. Their ages ranged from $15$ to $35$ years ($\text{MEAN}=26.625,\text{SD}=5.325$), comprising $13$ males and $3$ females. Participants included 1 high school student, 4 undergraduate students, 9 graduate students, and 2 research engineers. None of the participants had prior experience with AR/VR devices, which helped minimize bias due to varying familiarity with such technology.
@@ 4.1.2. Results Analysis
We evaluate GestureGPT’s performance by running the collected data through our system under four different context settings: 1) the baseline with only the function list (Baseline), 2) with gaze information (Only Gaze), 3) with interaction history and external information (Only History and External), and 4) all contexts are available (All). For comparison, we also provide results from Random Guess, which randomly selects one from all functions. We repeat the tests three times under each setting to ensure robust conclusions. The main results are presented in the left part of Table 2, and a corresponding illustration is provided in Figure 8.
@@
When comparing the benefits of gaze versus history and external information, we empirically find that gaze significantly outperforms history and external information. As shown in Table 2, the Only Gaze setting achieves 35.16%/83.59% at Top-1/Top-5 accuracy, outperforming the 23.18%/49.48% achieved by the Only History and External setting by a considerable margin. We attribute this to the semantic similarity between candidate functions shared by different home devices and their broad spatial distribution across the entire room space. These two factors significantly increase the reasoning difficulty under the Smart Home scenario. Facing such challenging situations, GestureGPT can fully exploit the gaze information as well as its own commonsense knowledge to finally locate and zoom in on possible candidates through multi-step, collaborative spatial reasoning. As explained in Section 3.3.1, when the Gesture Inference Agent requests gaze information, the Context Management Agent not only outputs the gaze coordinates but also identifies the relevant device within the gaze path using external tool-augmented Python scripts.
@@
Figure 8. Illustration of results from Experiment 1: Home Scenario, encompassing all 4 settings (across the x-axis) and 8 tasks (indicated by hue). For each setting, the 8 tasks are arranged from left to right as follows: task 1 Unlock Carbinet, task 2 Increase Light Brightness, task 3 Next Recipes, task 4 Open Oven, task 5 Open Air cleaner, task 6 Set Timer, task 7 Switch Input, and task 8 Phone Call. Each specific bar is further divided into Top 1 / Top 3 / Top 5 accuracy (Top 1, Top 3, Top 5 in gradient brown) and Negative outcomes (in blue), totaling 48 data points.
@@
Moreover, while integrating all context achieves the highest Top-1 performance, it unexpectedly underperforms in the Top-5 metric compared to the Gaze Only setting (79.69% vs 83.59%). Through further case analysis, we hypothesize that an excess of context can sometimes disrupt the agent’s analysis, leading to the following behaviors: 1) Irrelevant context information can at times mislead the agent. For example, the mention of “fingerprint unlocking” (in task 1) within external information might lead the agent to infer fingerprint recognition as the unlocking method, rather than a gesture. 2) An abundance of context causing the agent to overlook certain information. For example, in the “Open the air cleaner” task (in task 5) where the agent wrongly assumes the device is on, based on external information “The air purifier’s sensor detected heavy cooking fumes,” while ignoring its actual OFF status. These insights will be further discussed in Section 5.4 regarding optimal context selection.
@@
Finally, we observe that Tasks 2, 3, and 7 exhibit relatively high baseline performance, all involving operations related to “switching”. Finally, we observe that Tasks 2, 3, and 7 exhibit relatively high baseline performance, all involving operations related to “switching”. Similarly, for Tasks 3 and 7, GestureGPT employed a comparable analytical approach to discern the correct answer. This demonstrates GestureGPT’s strength in analyzing gestures alongside current device states, thereby accurately interpreting user intentions even in the baseline context.